1.  Consider the given reaction.

2BrF5(g)  ⇌ Br2(g) + 5F2(g)               

If the respective equilibrium concentrations of BrF5, Br2, and F2 at 250C are 0.046M, 1.2 M, and 0.93M, what are the values of Kc and Kp?


1
Expert's answer
2021-06-02T06:02:19-0400

Solution:

Balanced chemical equation:

2BrF5(g) ⇌ Br2(g) + 5F2(g)

The equilibrium expression (Kc) for the reaction:



Therefore,





We know that Kp = Kc (RT)Δn


Kc = 394.53

Convert 250C to Kelvin (K): T = 250C + 273.15 = 523.15 K

Δn = (total number of moles of products) - (total number of moles in reactants)

Δn = (5 + 1) - (2) = 6 - 2 = 4

Δn = 4

R = 0.08206 L atm mol-1 K-1


Now, plug in all the numbers we found: Kp = Kc (RT)Δn

Kp = 394.53 × (0.08206 × 523.15 K)4 = 1.34×109

Kp = 1.34×109


Answer:

Kc = 394.53

Kp = 1.34×109